OpenStudy (anonymous):

Which fraction is equivalent to 36%? A. 36/10 B.3 6/10 C.360/100 D.9/25

OpenStudy (alexandervonhumboldt2):

36/100

OpenStudy (alexandervonhumboldt2):

36/100=36/4/100/4=9/25

OpenStudy (anonymous):

okay thxs

OpenStudy (anonymous):

it cant be 36/100 because thats not an awser